# FedLoGe: Joint Local and Generic Federated Learning under Long-tailed Data

### 17 Jan 2024 v1

@@ Abstract
Federated Long-Tailed Learning (Fed-LT), a paradigm wherein data collected from decentralized local clients manifests a globally prevalent long-tailed distribution, has garnered considerable attention in recent times. In the context of Fed-LT, existing works have predominantly centered on addressing the data imbalance issue to enhance the efficacy of the generic global model while neglecting the performance at the local level. In contrast, conventional Personalized Federated Learning (pFL) techniques are primarily devised to optimize personalized local models under the presumption of a balanced global data distribution. This paper introduces an approach termed Federated Local and Generic Model Training in Fed-LT (FedLoGe), which enhances both local and generic model performance through the integration of representation learning and classifier alignment within a neural collapse framework. Our investigation reveals the feasibility of employing a shared backbone as a foundational framework for capturing overarching global trends, while concurrently employing individualized classifiers to encapsulate distinct refinements stemming from each client’s local features. Building upon this discovery, we establish the Static Sparse Equiangular Tight Frame Classifier (SSE-C), inspired by neural collapse principles that naturally prune extraneous noisy features and foster the acquisition of potent data representations. Furthermore, leveraging insights from imbalance neural collapse’s classifier norm patterns, we develop Global and Local Adaptive Feature Realignment (GLA-FR) via an auxiliary global classifier and personalized Euclidean norm transfer to align global features with client preferences. Extensive experimental results on CIFAR-10/100-LT, ImageNet, and iNaturalist demonstrate the advantage of our method over state-of-the-art pFL and Fed-LT approaches. Our codes are available at https://github.com/ZackZikaiXiao/FedLoGe.
@@ 1 Introduction
Nevertheless, existing works on Fed-LT have primarily focused on addressing the imbalance in the context of the global long-tailed data (Yang et al., 2023a; Qian et al., 2023), neglecting the tailoring of models to the needs of individual clients, since local data statistics could be diverse and not necessarily long-tailed. Personalized federated learning (pFL) (Tan et al., 2022), which trains customized local models for a single or a group of clients, offers an alternative solution to prioritize each client’s (or group’s) distinct data statistics and preferences, in which the global generic model is deemed as a bridge for training local personalized models and boosting the local performance with expressive representations (Li et al., 2021b; Collins et al., 2021; Li et al., 2021a). However, conventional pFL approaches are not supposed to attain a superior global generic model in Fed-LT.
@@
We examine the effectiveness of training with fixed classifiers in Fed-LT from the perspective of neural collapse (NC) (Papyan et al., 2020). The NC identifies a salient property of the feature space, showing that all within-same-class features tend to collapse to their respective class means. However, as shown in Fig. 1(a) [^1] Our $\psi_{SSE-C}$ leads to two notable improvements in Fig. 1 (b) over Fig. 1 (a): First, it masks noisy features with large variances, now in sparse areas marked by grey. Second, it enhanced dominant features, as shown by reduced variances among those with larger means, reflecting their increased precision and efficacy. , preliminary experiments benchmarking ETF with CIFAR-100 in Fed-LT suggest that only a few features have relatively large means, while most of the small-mean features are contaminated by severe noise. Such observations are inconsistent with the feature collapse property, and we coin it as feature degeneration. More details of the computation process for the data in Fig. 1 as well as the necessary explanation can be found in Appendix A.2.
@@ 2.1 Federated Learning
Personalized Federated Learning (pFL) To deal with the poor generalization performance of the single generic global at local data, pFL has been vastly investigated. A group of works seeks to train local personalized models via transferring knowledge from the generic global model (Li & Wang, 2019; Chen et al., 2020; T Dinh et al., 2020; Fallah et al., 2020; Chen et al., 2023). Multi-task learning-based methods are also been explored with client clustering (Sattler et al., 2020; Briggs et al., 2020; Ghosh et al., 2020) and model interpolation (Deng et al., 2020; Li et al., 2021a; Diao et al., 2020).
@@ 2.2 Neural Collapse for Representation Learning
Our method is the first to achieve personalization under the neural collapse framework, which can alleviate both the imbalanced data of the global model and provide personalization capabilities for the local models by adapting their feature distributions to their preferences simultaneously.
@@ 3.1 Preliminaries
where $\mathcal{D}=\{\mathcal{D}_{k}\}_{k=1}^{K}$ is the global long-tailed dataset composed of $K$ heterogeneous local datasets, each with size $|\mathcal{D}_{k}|$.
The training is executed over $T$ rounds. In each round $t$, the server distributes the current global model $w^{(t)}$ to all clients for local updates. Furthermore, given that $c$ represents the class index, for $\forall c\in C$, the classifier vector is $\mathbf{\psi}_{c}$, and the corresponding features are $h_{c}$.
@@ 3.2 Static Sparse Equiangular Tight Frame Classifier (SSE-C)
where $\mathbf{\psi}=[\mathbf{\psi}_{:,1},\cdots,\mathbf{\psi}_{:,C}]\in\mathbb{R}^{d\times C},\mathbf{U}\in\mathbb{R}^{d\times C}$ allows any rotation and satisfies $\mathbf{U}^{T}\mathbf{U}=\mathbf{I}_{C},\mathbf{I}_{C}$ is the identity matrix. $d$ is the dimension of the classifier vector, and $\mathbf{1}_{C}$ is an all-ones vector. We can deduce the important property that all class vector has the equal $\ell_{2}$ norm and maximal pair-wise angle $-\frac{1}{C-1}$ in $\mathbb{R}^{d}$. Note that randomly assigning $\beta$ proportion of weights in the ETF matrix to 00 will disrupt the ETF condition: the class vector angles cease to be maximal and equal, and the norms of the classifier vector become unequal. As such, it is necessary to train a sparse ETF structure that satisfies the ETF geometric conditions. We introduce a sparse indicator matrix $\mathbf{S}$ with the same dimensions as $\mathbf{\psi}$, where $\beta$ proportion of the elements are randomly set to 00. Then, the sparsified matrix $\mathbf{\psi}^{\prime}$ can be represented as $\mathbf{\psi}^{\prime}=\mathbf{\psi}\odot\mathbf{S}$. We design the Equal Norm Loss and Maximal Angle Loss to optimize the geometric structure to meet the conditions of ETF. First, the $\ell_{2}$ norm of class vectors should be equal. We constrain all norms of class vector values to a predetermined $\gamma$ with the formula:
@@ 4.1 Experimental Setup
Dataset, Models and Metrics:
We consider image classification tasks for performance evaluation on benchmark long-tailed datasets: CIFAR-10/100-LT, ImageNet-LT, and iNaturalist-User-160k (Van Horn et al., 2018). The CIFAR-10/100-LT datasets are sampled into a long-tailed distribution employing an exponential distribution governed by the Imbalance Factor (IF) in Cao et al. (2019). All experiments are conducted with non-IID data partitions, implemented by the Dirichlet distributions-based approach with parameter $\alpha$ to control the non-IIDness (Chen & Chao, 2022).ResNet-18 is trained over $K=40$ clients on CIFAR-10-LT, while ResNet-34 and ResNet-50 are implemented on CIFAR-100-LT and ImageNet-LT, respectively, with $K=20$ clients. The configurations for iNaturalist-160k align with those utilized for ImageNet-LT. We use $\alpha=1,0.5$ and IF = $50,100$ in CIFAR-10/100-LT. We use $\alpha=0.1,0.5$ in ImageNet-LT and iNaturalist, respectively.
@@
Compared Methods: In addition to FedAvg and FedProx (Li et al., 2020a) which are included for reference, we consider two types of state-of-the-art baselines: (1) pFL methods, including FedBN (Li et al., 2021b), FedPer (Arivazhagan et al., 2019), FedRep (Collins et al., 2021), Ditto (Li et al., 2021a), FedROD (Chen & Chao, 2022), have been applied during local training of each client. FedBN and FedRep incorporate all weights during aggregation, but exclude the batch normalization layers/classifiers during assignment, enabling the observation of the global model’s performance. (2) Federated (Long-tailed) Representation learning, including FedBABU (Oh et al., 2022), FedETF (Li et al., 2023) and Ratio Loss (Wang et al., 2021).
@@ 4.3 Ablation study and Sensitivity analysis
Effectiveness of SSE-C: We opted for a learning rate of 0.0001 with 10,000 optimization steps on ImageNet-LT with 1000 classes. The order of magnitude for the norm variance is between $10^{-7}$ and $10^{-9}$. For the angles across each pair of classifier vectors, we visualized the cosine similarity in SSE-C in Fig. 3, indicating that our optimized SSE-C is almost identical to the dense ETF.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 3: (a): The heatmap of cosine similarity between classifier vectors in SSE-C; (b): Pruning features from smaller means and larger means, respectively. Negligible features exert a minor impact on model performance; (c): Pruning Experiments with SSE-C. The model optimally enhances the dominant features, rendering the impact of negligible features imperceptible.
@@
Negligible and dominant features with SSE-C: We investigated the effects of pruning on different features over CIFAR-100-LT ($\alpha$=0.5, IF=100). We computed the class mean for each category. Within each class mean, we ordered the means and pruned the corresponding classifier vector weights in descending (from dominant features) and ascending order (from negligible features) with pruning ratios from 0 to $100\%$. After training with SSE-C, pruning dominant features drastically decreases performance, while pruning negligible features barely affects performance, which is visualized in Fig. 3 (b). This observation indicates that SSE-C can adeptly learn more effective features while autonomously disregarding high-noise features.
@@ A.1 Detailed setup
A.2 Detailes of feature degeneration
Given that $n_{c}$ is the sample number of class $c$. $\mu_{c}$ is the class mean of class $c$, and $h_{i,c}$ is the feature of class $c$ on sample $i$, the features $h_{i,c}$ will collapse to the within-class mean $\mu_{c}=\frac{1}{n_{c}}\sum_{i=1}^{n_{c}}h_{i,c}$. This indicates that the covariance $\frac{1}{n_{c}}\sum_{i=1}^{n_{c}}(h_{i,c}-\mu_{c})(h_{i,c}-\mu_{c})^{T}$ will converge to 00. We investigate whether $h_{i,c}-\mu_{c}$ tends toward zero when training with a fixed ETF classifier, as shown in Fig.  1. The steps to obtain the data in Fig.  1 are: 1. Select class $c$. 2. During global test, compute all sample features $h_{i,c}$ with dimension $D$, representing the number of features. Use $h_{i,c}[d]$ to denote the $d^{th}$ feature in $h_{i,c}$. 3. Compute the mean and relative variance of $h_{c}$, sorting them in descending order of the mean, and accordingly adjust the variance positions. It becomes evident that not all features, $h_{i,c}[d]-\mu_{c}[d]$, converge to zero, and smaller $\mu_{c}[d]$ values exhibit greater convergence noise. The process illustrated in Fig.  1 (b) follows the same steps as Fig.  1 (a), but with grey vertical shadows for the sparsely positioned features. It can be observed that sparse ETF training effectively squeezes poor-quality and noisy features into these sparsely allocated positions.
A.3 Computational Cost
@@
A.4 Local Finetune Further Improve Performance
@@
A.5 T-SNE Visualization of Class Means
@@ A.6 Norm and Angles of SSE-C
A.7 Local Norm or local distribution for LA-FR
@@
On CIFAR100-LT with $\alpha=0.5$ and $\text{IF}=100$, we realign personalized models employing both norm and distribution, calculating the arithmetic mean, macro average, and weighted average of all clients’ performance. The results can be seen in Tab.  6. We discern that utilizing norm for realignment is more effective than employing the real distribution. One rationale is that the norm more aptly signifies the model’s cognitive capacity over the dataset. For instance, although class A encompasses more samples than class B, B is more prone to misprediction owing to its high resemblance with other classes. In this scenario, realignment ought to adhere to model cognition, fortifying class A over B.
@@
A.8 Projection Layer and Dot-regression Loss is Not Necessary for Fed-LoGe

### 08 Mar 2024 v2

@@ Abstract
Federated Long-Tailed Learning (Fed-LT), a paradigm wherein data collected from decentralized local clients manifests a globally prevalent long-tailed distribution, has garnered considerable attention in recent times. In the context of Fed-LT, existing works have predominantly centered on addressing the data imbalance issue to enhance the efficacy of the generic global model while neglecting the performance at the local level. In contrast, conventional Personalized Federated Learning (pFL) techniques are primarily devised to optimize personalized local models under the presumption of a balanced global data distribution. This paper introduces an approach termed Federated Local and Generic Model Training in Fed-LT (FedLoGe), which enhances both local and generic model performance through the integration of representation learning and classifier alignment within a neural collapse framework. Our investigation reveals the feasibility of employing a shared backbone as a foundational framework for capturing overarching global trends, while concurrently employing individualized classifiers to encapsulate distinct refinements stemming from each client’s local features. Building upon this discovery, we establish the Static Sparse Equiangular Tight Frame Classifier (SSE-C), inspired by neural collapse principles that naturally prune extraneous noisy features and foster the acquisition of potent data representations. Furthermore, leveraging insights from imbalance neural collapse’s classifier norm patterns, we develop Global and Local Adaptive Feature Realignment (GLA-FR) via an auxiliary global classifier and personalized Euclidean norm transfer to align global features with client preferences. Extensive experimental results on CIFAR-10/100-LT, ImageNet-LT, and iNaturalist demonstrate the advantage of our method over state-of-the-art pFL and Fed-LT approaches. Our codes are available at https://github.com/ZackZikaiXiao/FedLoGe.
@@ 1 Introduction
A high-quality global model with balanced performance could speed up local adaptation and attract new clients, while personalized models aim to provide enhanced local performance by considering local data characteristics. Nevertheless, existing works on Fed-LT have primarily focused on addressing the imbalance in the context of the global long-tailed data (Yang et al., 2023a; Qian et al., 2023; Xiao et al., 2023), neglecting the tailoring of models to the needs of individual clients, since local data statistics could be diverse and not necessarily long-tailed. Personalized federated learning (pFL) (Tan et al., 2022), which trains customized local models for a single or a group of clients, offers an alternative solution to prioritize each client’s (or group’s) distinct data statistics and preferences, in which the global generic model is deemed as a bridge for training local personalized models and boosting the local performance with expressive representations (Li et al., 2021b; Collins et al., 2021; Li et al., 2021a). However, conventional pFL approaches are not supposed to attain a superior global generic model in Fed-LT.
@@
We examine the effectiveness of training with fixed classifiers in Fed-LT from the perspective of neural collapse (NC) (Papyan et al., 2020). The NC identifies a salient property of the feature space, showing that all within-same-class features tend to collapse to their respective class means. However, as shown in Fig. 1(a) [^1] Our $\psi_{SSE-C}$ leads to two notable improvements in Fig. 1 (b) over Fig. 1 (a): First, it masks noisy features with large variances, now in sparse areas marked by grey. Second, the role of dominant features is enhanced, as shown by reduced variances among those with larger means, reflecting their increased precision and efficacy. , preliminary experiments benchmarking ETF with CIFAR-100 in Fed-LT suggest that only a few features have relatively large means, while most of the small-mean features are contaminated by severe noise. Such observations are inconsistent with the feature collapse property, and we coin it as feature degeneration. More details of the computation process for the data in Fig. 1 as well as the necessary explanation can be found in Appendix A.2.
@@ 2.1 Federated Learning
Personalized Federated Learning (pFL) To deal with the poor generalization performance of the single generic global at local data, pFL has been vastly investigated. A group of works seeks to train local personalized models via transferring knowledge from the generic global model (Li & Wang, 2019; T Dinh et al., 2020; Fallah et al., 2020; Chen et al., 2023). Multi-task learning-based methods are also been explored with client clustering (Sattler et al., 2020; Briggs et al., 2020; Ghosh et al., 2020) and model interpolation (Deng et al., 2020; Li et al., 2021a; Diao et al., 2020).
@@ 2.2 Neural Collapse for Representation Learning
In summary, our method stands out by introducing personalization within the neural collapse framework, effectively overcoming global imbalanced data and enhancing local model personalization through tailored feature distribution alignment. Contrary to FedETF’s fixed classifier approach and FedRod’s dual classifier structure, our method ensures superior handling of global long-tail bias and precise local model tuning, marking a novel advancement in Fed-LT. For detailed comparisons, please see Section A.13 in the Appendix.
@@ 3.1 Preliminaries
where $\mathcal{D}=\{\mathcal{D}_{k}\}_{k=1}^{K}$ is the global long-tailed dataset composed of $K$ heterogeneous local datasets, each with size $|\mathcal{D}_{k}|$. The training is executed over $T$ rounds. In each round $t$, the server distributes the current global model $w^{(t)}$ to all clients for local updates. Furthermore, given that $c$ represents the class index, for $\forall c\in C$, the classifier vector is $\mathbf{\psi}_{c}$, and the corresponding features are $h_{c}$.
@@ 3.2 Static Sparse Equiangular Tight Frame Classifier (SSE-C)
where $\mathbf{\psi}=[\mathbf{\psi}_{:,1},\cdots,\mathbf{\psi}_{:,C}]\in\mathbb{R}^{d\times C},\mathbf{U}\in\mathbb{R}^{d\times C}$ allows any rotation and satisfies $\mathbf{U}^{T}\mathbf{U}=\mathbf{I}_{C},\mathbf{I}_{C}$ is the identity matrix. $d$ is the dimension of the classifier vector, and $\mathbf{1}_{C}$ is an all-ones vector. We can deduce the important property that all class vector has the equal $\ell_{2}$ norm and maximal pair-wise angle $-\frac{1}{C-1}$ in $\mathbb{R}^{d}$. Note that randomly assigning $\beta$ proportion of weights in the ETF matrix to 00 will disrupt the ETF condition: the class vector angles cease to be maximal and equal, and the norms of the classifier vector become unequal. As such, it is necessary to train a sparse ETF structure that satisfies the ETF geometric conditions. We introduce a sparse indicator matrix $\mathbf{S}$ with the same dimensions as $\mathbf{\psi}$, where $\beta$ proportion of the elements are randomly set to 00. Then, the sparsified matrix $\mathbf{\psi}^{\prime}$ can be represented as $\mathbf{\psi}^{\prime}=\mathbf{\psi}\odot\mathbf{S}$. We design the Equal Norm Loss and Maximal Angle Loss to optimize the geometric structure to meet the conditions of ETF. First, the $\ell_{2}$ norm of class vectors should be equal. We constrain all norms of class vector values to a predetermined $\gamma$:
@@ 4.1 Experimental Setup
Dataset, Models and Metrics: We consider image classification tasks for performance evaluation on benchmark long-tailed datasets: CIFAR-10/100-LT, ImageNet-LT, and iNaturalist-User-160k (Van Horn et al., 2018). The CIFAR-10/100-LT datasets are sampled into a long-tailed distribution employing an exponential distribution governed by the Imbalance Factor (IF) in Cao et al. (2019). All experiments are conducted with non-IID data partitions, implemented by the Dirichlet distributions-based approach with parameter $\alpha$ to control the non-IIDness (Chen & Chao, 2022).ResNet-18 is trained over $K=40$ clients on CIFAR-10-LT, while ResNet-34 and ResNet-50 are implemented on CIFAR-100-LT and ImageNet-LT, respectively, with $K=20$ clients. The configurations for iNaturalist-160k align with those utilized for ImageNet-LT. We use $\alpha=1,0.5$ and IF = $50,100$ in CIFAR-10/100-LT. We use $\alpha=0.1,0.5$ in ImageNet-LT and iNaturalist, respectively.
@@
Compared Methods: In addition to FedAvg and FedProx (Li et al., 2020a) which are included for reference, we consider two types of state-of-the-art baselines: (1) pFL methods, including FedBN (Li et al., 2021b), FedPer (Arivazhagan et al., 2019), FedRep (Collins et al., 2021), Ditto (Li et al., 2021a), and FedROD (Chen & Chao, 2022). (2) Federated (Long-tailed) Representation learning, including FedBABU (Oh et al., 2022), FedETF (Li et al., 2023) and Ratio Loss (Wang et al., 2021).
@@ 4.3 Ablation study and Sensitivity analysis
Refer to caption Refer to caption
(a) (b)
Figure 3: (a): Pruning features from smaller means and larger means, respectively. Negligible features exert a minor impact on model performance; (b): Pruning Experiments with SSE-C. The model optimally enhances the dominant features, rendering the impact of negligible features imperceptible.
Negligible and dominant features with SSE-C: We investigated the effects of pruning on different features over CIFAR-100-LT ($\alpha$=0.5, IF=100). We computed the class mean for each category. Within each class mean, we ordered the means and pruned the corresponding classifier vector weights in descending (from dominant features) and ascending order (from negligible features) with pruning ratios from 0 to $100\%$. After training with SSE-C, pruning dominant features drastically decreases performance, while pruning negligible features barely affects performance, which is visualized in Fig. 3 (a, b). This observation indicates that SSE-C can adeptly learn more effective features while autonomously disregarding high-noise features.
@@ 5 Conclusion
Acknowledgements
This work is supported by the National Natural Science Foundation of China (Grant No. 62106222, No. 62201504), the Natural Science Foundation of Zhejiang Province, China(Grant No. LZ23F020008, No. LGJ22F010001), Zhejiang Lab Open Research Project (No. K2022PD0AB05) and the Zhejiang University-Angelalign Inc. R&\&&D Center for Intelligent Healthcare.
@@ A.1 Detailed setup
A.2 Table of Notations
Please refer to Table 4 for the table of notations used throughout this paper.
Notation Description
$K$ Number of clients in the federated learning system.
$w$ Parameters of the generic model, consisting of $\theta$ and $\mathbf{\psi}$.
$w_{k}$ Parameters of the $k^{\text{th}}$ personalized model for client $k$.
$f(x,\theta)$ Feature extractor parameterized by $\theta$.
$h$ Features transformed from input $x$ by feature extractor $f$.
$g(h,\mathbf{\psi})$ Generic classifier mapping features $h$ to output labels.
$g(h,\mathbf{\phi}_{k})$ Personalized classifier for client $k$.
$\mathbf{\phi}_{k}$ Personalized classifier parameters for client $k$.
$\mathcal{D}$ Global long-tailed dataset.
$\mathcal{D}_{k}$ Local dataset of client $k$.
$T$ Number of training rounds.
$c$ Class index.
$\mathbf{\psi}_{c}$ Classifier vector corresponding to class $c$.
$h_{c}$ Features corresponding to class $c$.
$\mathbf{S}$ Sparse indicator matrix.
$\gamma$ Predetermined $\ell_{2}$ norm value for class vectors.
$\mathcal{L}_{\text{SSE-C}}$ Loss function for Static Sparse Equiangular Tight Frame Classifier (SSE-C).
$\mathcal{L}_{\text{norm}}$, $\mathcal{L}_{\text{angle}}$ Loss functions for optimizing the geometric structure of SSE-C.
$n_{c}$ Sample number of class $c$.
$\mu_{c}$ Class mean of features for class $c$.
$h_{i,c}$ Feature of class $c$ for sample $i$.
Table 4: Summary of notations used in the paper.
A.3 Detailes of feature degeneration
Given that $n_{c}$ is the sample number of class $c$. $\mu_{c}$ is the class mean of class $c$, and $h_{i,c}$ is the feature of class $c$ on sample $i$, the features $h_{i,c}$ will collapse to the within-class mean $\mu_{c}=\frac{1}{n_{c}}\sum_{i=1}^{n_{c}}h_{i,c}$. This indicates that the covariance $\frac{1}{n_{c}}\sum_{i=1}^{n_{c}}(h_{i,c}-\mu_{c})(h_{i,c}-\mu_{c})^{T}$ will converge to 00. We investigate whether $h_{i,c}-\mu_{c}$ tends toward zero when training with a fixed ETF classifier, as shown in Fig.  1. The steps to obtain the data in Fig.  1 are: 1. Select class $c$. 2. During global test, compute all sample features $h_{i,c}$, representing the number of features. Use $h_{i,c}[d]$ to denote the $d^{th}$ feature in $h_{i,c}$. 3. Compute the mean and relative variance of $h_{c}$, sorting them in descending order of the mean, and accordingly adjust the variance positions. It becomes evident that not all features, $h_{i,c}[d]-\mu_{c}[d]$, converge to zero, and smaller $\mu_{c}[d]$ values exhibit greater convergence noise. The process illustrated in Fig.  1 (b) follows the same steps as Fig.  1 (a), but with grey vertical shadows for the sparsely positioned features. It can be observed that sparse ETF training effectively squeezes poor-quality and noisy features into these sparsely allocated positions.
A.4 Computational Cost
@@
A.5 Local Finetune Further Improve Performance
@@
A.6 T-SNE Visualization of Class Means
@@ A.7 Norm and Angles of SSE-C
A.8 Local Norm or local distribution for LA-FR
@@
On CIFAR100-LT with $\alpha=0.5$ and $\text{IF}=100$, we realign personalized models employing both norm and distribution, calculating the arithmetic mean, macro average, and weighted average of all clients’ performance. The results can be seen in Tab.  7. We discern that utilizing norm for realignment is more effective than employing the real distribution. One rationale is that the norm more aptly signifies the model’s cognitive capacity over the dataset. For instance, although class A encompasses more samples than class B, B is more prone to misprediction owing to its high resemblance with other classes. In this scenario, realignment ought to adhere to model cognition, fortifying class A over B.
@@
A.9 Projection Layer and Dot-regression Loss is Not Necessary for Fed-LoGe